# A Backbone for Long-Horizon Robot Task Understanding

### v1

@@ III-C ActionREG: SAM-driven action registration network
Through ActionREG, we can easily extract the task-related object mask and the operational workspace mask. Utilizing a hierarchical matching system based on SAM (Segment Anything Model), YOLOv8, SIFT (Scale-Invariant Feature Transform), FLANN (Fast Library for Approximate Nearest Neighbors), and PCA (Principal Component Analysis), we can identify task-related objects and determine their configurations in new scenarios. Specifically, $\mathbf{M}=\bigcup_{k\in\mathbf{K}}M_{\text{SAM}}(\mathbf{H}\cdot\mathcal{G}(\mathbf{S}_{\text{therbligs}},k),\mathbf{B})$ represents object mask segmentation using the SAM model, while $\mathbf{Box}_{k}=M_{\text{YOLO}}(\mathbf{I}_{\text{new}},\text{ComputeArea}(\mathbf{m}_{k}))$ denotes bounding box detection through YOLOv8. The feature extraction is handled by $\mathbf{F}_{\text{new}}=\text{SIFT}(\mathbf{Box}_{k})$ using SIFT, feature matching is performed by $\mathbf{M}_{\text{match}}=\text{FLANN}(\mathbf{F}_{\text{new}},\mathbf{F}_{\text{ref}})$ using FLANN, and orientation extraction is achieved via $\mathbf{O}_{\text{new},k}=\text{PCA}(\mathbf{M}_{\text{match}})$ using PCA.
@@ V Results and Analysis
Figure 7: Therblig segmentation recall for different robot tasks. It shows the recall rates for various therbligs across multiple robotic tasks, highlighting the performance of our segmentation network.
Furthermore, as shown in Fig. 8, the LAP-VC system consistently achieves high alignment performance scores across various tasks, outperforming traditional methods like KNN, SIFT, ORB, AKAZE, FAST, and BRISK. In particular, for the Roller task, the LAP-VC system achieved a score of 0.92, which is second only to the manual alignment score of 0.98. Similarly, in the Spoon task, the LAP-VC scored 0.88, close to the perfect manual score of 1.00. In the Stamp, Sponge, and Scraper tasks, the LAP-VC system scored 0.84, 0.94, and 0.90 respectively, showing a consistent high performance. Compared to other automated methods, LAP-VC shows superior robustness and reliability.

### v2

@@ III-C ActionREG: SAM-driven action registration network
Through ActionREG, we can easily extract the task-related object mask and the operational workspace mask. Utilizing a hierarchical matching system based on SAM (Segment Anything Model), YOLOv8 (You only look once version 8), SIFT (Scale-Invariant Feature Transform), FLANN (Fast Library for Approximate Nearest Neighbors), and PCA (Principal Component Analysis), we can identify task-related objects and determine their configurations in new scenarios. Specifically, $\mathbf{M}=\bigcup_{k\in\mathbf{K}}M_{\text{SAM}}(\mathbf{H}\cdot\mathcal{G}(\mathbf{S}_{\text{therbligs}},k),\mathbf{B})$ represents object mask segmentation using the SAM model, while $\mathbf{Box}_{k}=M_{\text{YOLO}}(\mathbf{I}_{\text{new}},\text{ComputeArea}(\mathbf{m}_{k}))$ denotes bounding box detection through YOLOv8. The feature extraction is handled by $\mathbf{F}_{\text{new}}=\text{SIFT}(\mathbf{Box}_{k})$ using SIFT, feature matching is performed by $\mathbf{M}_{\text{match}}=\text{FLANN}(\mathbf{F}_{\text{new}},\mathbf{F}_{\text{ref}})$ using FLANN, and orientation extraction is achieved via $\mathbf{O}_{\text{new},k}=\text{PCA}(\mathbf{M}_{\text{match}})$ using PCA.
@@ V Results and Analysis
Figure 7: Therblig segmentation recall for different robot tasks. It shows the recall rates for various therbligs across multiple robotic tasks, highlighting the performance of our segmentation network. For picka and place task, data points represented with an underscore indicate that the data is none with use.
Furthermore, as shown in Fig. 8, the LAP-VC system consistently achieves high alignment performance scores, averaging 0.896 across various tasks, outperforming traditional methods such as KNN, SIFT, ORB, AKAZE, FAST, and BRISK. Its performance is only slightly lower than that of human experts manually performing alignment. Specifically, for the Stamp task, the LAP-VC system achieved a score of 0.84. This lower score is attributed to the relatively small size of the stamp. In contrast, the Sponge task achieved a higher score of 0.94 due to the sponge’s relatively simple and uniform structure.